OpenStudy (anonymous):

Write a linear factorization of the function. f(x) = x4+ 9x2 f(x) = x2(x + 3i)(x - 3i) f(x) = x2(x + 3i)2 f(x) = x2(3x + i)(3x - i) f(x) = x2(3x + i)

OpenStudy (anonymous):

f(x) = x2(3x + i)^2

OpenStudy (inkyvoyd):

Here's a hint: (a^2+b^2)=(a+bi)(a-bi) hint hint.

OpenStudy (anonymous):

(3x+i)(3x-i)

OpenStudy (inkyvoyd):

Nope.

OpenStudy (inkyvoyd):

Lemme rewrite it. \(\Huge (x^{2})^{2}+(3x)^2\) hinty hint.

OpenStudy (anonymous):

i dont understand @inkyvoyd how can you get an answer i think its f(x) = x2(3x + i)(3x - i)

OpenStudy (anonymous):

f(x) = x2(x + 3i)2 i think this @inkyvoyd

OpenStudy (inkyvoyd):

Well \(\Huge (x^{2})^{2}+(3x)^2\) \(\huge\text{a=}x^2\text{b=}3x\)

OpenStudy (inkyvoyd):

\(\Huge a^2+b^2=(a+bi)(a-bi)\)

OpenStudy (anonymous):

f(x) = x2(x + 3i)(x - 3i) ohhhhhhhhhhh @inkyvoyd lol

OpenStudy (inkyvoyd):

Yup.